OpenStudy (anonymous):

There are two factors of -36 such that one factor is 11 less than half of the other factor. Choose all the pairs of these factors. -2 and 18 -6 and 6 3 and -12 4 and -9

OpenStudy (anonymous):

yikes lets try this one "-2 and 18" half of 18 is 9, and \(9-11=-2\) so i guess we got it on the first try

OpenStudy (anonymous):

is there more than one?

OpenStudy (anonymous):

yeah i think the last one might work

OpenStudy (anonymous):

half of 4 is 2, and \(2-11=-9\) maybe there is another as well

OpenStudy (anonymous):

half of 6 is 3, and \(3-11=-8\) so 6,-6 doesn't work

OpenStudy (anonymous):

and half of three isn't even an integer so that one is not going to work either

OpenStudy (anonymous):

oaky thank you so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

okay*

OpenStudy (anonymous):

otay